If a pharmacist adds 6 g of pure hydrocortisone to 97 g of an ointment which contains 5 % w/w hydrocortisone, what is the final concentration (% w/w) of hydrocortisone in the product? (1 decimal place)
Answer #1

The given ointment has 5% w/w hydrocortisone.

weight of ointment is 97 grams

weight of hydrocortisone in 97 grams of ointment is = (%w/w of hydrocortisone * weight of ointment) / 100

which is (5*97)/100 = 4.85 grams

4.85 grams of hydrocortisone present in 97 grams of ointment.

weight of pure hydrocortisone added is 6 grams.

Total weight of hydrocortisone in ointment after addition is 6 + 4.85 = 10.85 grams

Final weight of ointment is = previous weight + weight or hydrocortisone added

which is 97 + 6 = 103.

w/w concentration of hydrocortisone after addition is =

( weight of hydrocortisone after addition * 100 )/ total final weight of ointment

which is (10.85 * 100)/103 = 10.53 % w/w which on rounding to one decimal point becomes 10.5

The final concentration of hydrocortisone is 10.5 % w/w

Hope this helps you. Please feel free to give a like and contact us if you need further detailed explanation. Thank you and have a great day.